# ArXiv Preprint

### v1

@@ Abstract
Recent advances in large language models (LLMs) enable compelling story generation, but connecting narrative text to playable visual environments remains an open challenge in procedural content generation (PCG). We present a lightweight pipeline that transforms short narrative prompts into a sequence of 2D tile-based game scenes, reflecting the temporal structure of stories. Given an LLM-generated narrative, our system identifies three key time frames, extracts spatial predicates in the form of "Object-Relation-Object" triples, and retrieves visual assets using affordance-aware semantic embeddings from the GameTileNet dataset [1]. A layered terrain is generated using Cellular Automata, and objects are placed using spatial rules grounded in the predicate structure. We evaluated our system in ten diverse stories, analyzing tile–object matching, affordance–layer alignment, and spatial constraint satisfaction across frames. This prototype offers a scalable approach to narrative-driven scene generation and lays the foundation for future work on multi-frame continuity, symbolic tracking, and multi-agent coordination in story-centered PCG.
@@ 1 Introduction
Large language models (LLMs) have advanced the capacity to generate rich, coherent narratives from minimal input. This progress opens new possibilities for procedural content generation (PCG), especially in games that rely on narrative depth and visual storytelling. However, while narrative generation has improved significantly, the translation of these narratives into structured, spatially grounded game content remains an open challenge. Existing PCG systems often focus on producing isolated levels or static layouts, without addressing the continuity and temporal structure that characterize most narrative experiences.
Narratives unfold across time, comprising distinct events and evolving contexts that contribute to spatial progression, character development, and thematic coherence [2, 3]. Motivated by this structure, we propose a method for generating sequences of visual game scenes based on LLM-generated narratives. Rather than visualizing a single static moment, our system segments each story into three time frames and renders them as layered, tile-based game scenes. This enables partial modeling of narrative progression and sets the stage for future work on temporal consistency and dynamic storytelling in visual game environments.
To bridge narrative language and spatial visual representation, we design a lightweight pipeline that extracts symbolic spatial predicates from each time frame, matches them with visual assets from the GameTileNet dataset [1], and arranges them on procedurally generated terrains. Using "Object-Relation-Object" triples derived through LLM prompting, each object is semantically matched to one or more game tiles, constrained by their labeled affordances. Terrain layers are generated using Cellular Automata to ensure spatial connectivity, and objects are placed according to inferred spatial relations. The resulting scenes are layered 2D renderings that reflect the structure and intent of the source narrative while preserving game-relevant spatial logic.
Table 1: An example of generating a tile-based game scene from narrative descriptions. The system extracts spatial relations and semantic objects (left), selects appropriate tile images from the GameTileNet dataset (middle), and renders a visual scene that satisfies both semantic and spatial constraints (right). Top image: labeled object layout. Bottom image: tile-mapped game scene.
@@
We evaluate the system using ten narrative examples, each divided into three time frames. We assess the quality of tile-object matching, alignment between object affordances and scene layers, and the degree to which spatial relations are satisfied in the final renderings. Results show that even with lightweight symbolic rules and open-ended input, the generated scenes maintain semantic coherence and reflect narrative structure.
This work contributes the following:
  • A pipeline that decomposes LLM-generated narratives into temporally segmented scenes, rendered as layered 2D game environments.
  • A predicate extraction and semantic-matching strategy that grounds narrative objects in tile-based visual representations, using affordance-aligned filtering.
  • A multi-layer scene synthesis method that integrates procedurally generated terrain with symbolic spatial placement rules derived from narrative predicates.
  • An analysis of ten narrative examples showing semantic alignment, affordance-layer coherence, and spatial relation satisfaction across time frames.
Our system presents a modular framework for narrative-to-scene generation, emphasizing temporal segmentation, affordance grounding, and semantic alignment. While the current prototype does not model explicit state transitions or agent behaviors, it offers a foundation for future work on temporal continuity, character persistence, and multi-agent coordination in narrative-driven PCG.
Recent findings suggest that LLMs are capable of segmenting narrative events in ways that align with human judgments [4], supporting our approach of extracting structured time frames as a basis for visual scene construction. Table 1 illustrates one example of the output, highlighting how spatial constraints and tile affordances together shape semantically coherent game scenes from narrative descriptions.
@@ 2.1 Narrative Theory and Structure in Games
Narrative structure in games has long been explored through diverse lenses, from formalist definitions of interactivity to emergent storytelling models. Early foundational works examined how game rules intersect with narrative progression, emphasizing the tension between linear storytelling and player agency [5, 6, 7]. Building on this, various models have been proposed to categorize narrative mechanisms, such as experiential storytelling [8], story architectures that accommodate player agency [9], and frameworks to avoid narrative-schema conflict in game design [10].
Complementary perspectives have further connected narrative to spatial and ludic design elements. For example, narrative logic has been analyzed in terms of spatial layout and symbolic environments [11, 12], while encyclopedic works have surveyed narrative patterns in interactive media [13, 14]. Empirical studies of interactive storytelling also emphasize player engagement and agency [15, 16]. These theories form the basis for incorporating narrative intent into procedural generation systems.
@@ 2.2 Narrative-Guided Procedural Content Generation
Recent work has explored how narrative structure can guide procedural content generation (PCG), particularly in level and quest generation. One line of research focuses on generating quests or narrative arcs using language models or symbolic planning systems. For instance, story grammars and quest ontologies have been used to structure goal-driven content [17, 18], while large language models have been adapted to generate game worlds, characters, and events [19, 20, 21, 22]. Other work integrates emotional arcs or planning constraints to improve coherence and player engagement [23, 24, 25].
Several frameworks take a hybrid approach, combining symbolic control with neural methods. Procedural generation by content type has been classified according to the use of generative grammars, search-based methods, and PCGML [26, 27, 28]. More recently, reinforcement learning and evolutionary methods have been applied to PCG systems for levels or narrative structuring [29, 30].
Layered generation approaches have also emerged in the domain of visual narrative generation, including in 2D comics. Previous methods have used hierarchical rendering pipelines that maintain semantic control across visual layers (e.g., layout, background, foreground, characters), offering control over narrative-to-visual mappings. Such work demonstrates the feasibility of scene composition from structured narrative input, though not yet fully aligned with game semantics [31, 32].
@@ 2.3 Tile-based generation and semantic affordances
Procedural layout generation often leverages tile-based abstractions for scalability and modularity. Classical approaches include constrained layout generation via SMT solvers [33], evolutionary or multi-objective search for scene composition [34], and grammar-based or dynamic convolution systems for indoor or dungeon layouts [35, 36]. More recent neural approaches propose tile embeddings for generalized level generation [37] or semantic segmentation systems to produce context-aware tilemaps [38].
Semantic affordance labeling has been used to enable gameplay-aware modeling of objects and scenes. Prior taxonomies distinguish environmental objects, interactive elements, and collectibles [39, 40], while tile corpora and datasets have emerged to support affordance-aware generation [41]. Several frameworks combine rule-based labeling, pre-defined vocabularies (e.g., VGDL), and annotations of character-object interactions to enhance game-like semantics. The GameTileNet dataset contributes to this line by introducing a low-resolution game tile dataset with structured annotations and layered object affordance labels [1].
@@ 2.4 Semantic Matching and LLM-Guided Grounding
Large language models (LLMs) are increasingly employed in scene synthesis, semantic alignment, and grounding tasks. Recent work shows how prompting techniques and vision-language grounding can support scene and story generation, enabling LLMs to function as layout planners or tile selectors [42, 43]. Research also highlights how multimodal LLMs benefit from structured input, such as scene semantics, object descriptions, or event-level guidance [44, 45].
Visual-semantic reasoning frameworks further explore alignment between images and textual descriptions. These include methods for image-text matching, semantic embedding, and reasoning disentangled from raw perceptual input [46, 47, 48]. Recent cognitive studies suggest that LLMs may even approximate human-like segmentation of narrative events, underscoring their promise in symbolic-to-neural mapping tasks [4].
@@ 2.5 Symbolic Structures in Visual Storytelling
Graph-based or symbolic approaches to storytelling aim to encode structure and logic in narrative generation systems. Knowledge-enhanced storytelling has been explored using structured data, narrative graphs, or multimodal scene graphs [49, 50, 51, 52]. Frameworks like Aesop [meo2019aesop] or hierarchical story encoders [53] focus on visual or conversational storytelling with symbolic inputs.
Other work highlights how symbolic representations support event reasoning, narrative coherence, or compositional generation. For example, relational planning or temporal segmentation can support continuity and causality in game environments [54, 55]. Hierarchical narrative graph representations have also been proposed to support multimodal alignment, from sentence structure to visual frames in comics [56, 57]. Together, these systems underscore the benefit of symbolic scaffolds in guiding controllable, interpretable storytelling pipelines.
Together, these bodies of work provide a multi-aspect foundation for our approach. Theories of narrative structure and procedural generation highlight the challenges of preserving coherence and player agency in generative systems. Tile-based generation and affordance modeling offer concrete methods for grounding abstract narrative elements in spatialized, game-relevant formats. Recent progress in semantic matching and alignment between vision and language, particularly with large language models, suggests viable techniques for bridging textual story predicates with visual tile selection. Finally, symbolic representations in narrative and visual storytelling support the need for structured, interpretable intermediaries to mediate between story-driven goals and multi-layered game scene realization. Our proposed pipeline builds on and connects these strands, with the aim of advancing procedural generation based on narrative through structured, affordance-aware, and semantically aligned design.
@@ 3 Method: Narrative-to-Scene Generation Pipeline
We propose a structured pipeline that transforms narrative text into visually grounded, tile-based game scenes. As illustrated in Figure 1, the pipeline integrates large language model (LLM)–based story generation with a sequence of symbolic and visual reasoning modules. The process begins with narrative prompting and temporal abstraction, where the story is segmented into key time frames and represented as predicate-style triples. These structured descriptions serve as input for semantic reasoning and symbolic grounding.
Next, each scene is synthesized through a layered generation process: (1) terrain generation using Cellular Automata ensures navigable base regions; (2) semantic object matching retrieves tile assets aligned with narrative entities based on name, category, and affordance embeddings; and (3) spatial placement and rule-based refinement position entities according to their annotated spatial relations. The final output consists of rendered tile-based scenes and underlying semantic maps, enabling both human-readable visualization and downstream interactive applications.
Figure 1: Pipeline for generating a tile-based game scene from LLM-generated narrative. Modules include narrative parsing, semantic matching, procedural terrain generation, and layered scene rendering.
@@ 3.1 Narrative Prompting and Frame Extraction
Inspired by Word2World [19], we use a large language model (LLM) to generate a short adventure story and extract three key narrative moments. Each moment is then structured as a set of symbolic predicates that describe relationships between story elements. We refer to these as triples "object–relation–object", formally represented in the form [Object] [Relation] [Object]. These structured predicates form the basis for downstream spatial reasoning and the generation of visual scenes.
The process involves two prompts:
  • Prompt_1: "Generate a short adventure story (about 100 words)."
  • Prompt_2: "Extract three key time frames and describe each with [Object] [Relation] [Object] triplets."
@@
Table 2 illustrates the extracted time frames, the corresponding [Object] [Relation] [Object] predicates, and the tile-based scenes rendered by our pipeline.
@@
Hollow oak contains
ancient map
Elara stands near
hollow oak
Sunlight filters through
forest canopy
Elara walks along
rocky path
Wild creatures hide behind
dense bushes
Treacherous paths lead to
Crystal Cavern
Crystal Cavern entrance glows with
shimmering light
Guardian dragon sits atop
crystal throne
Elara stands before
Guardian dragon
@@ 3.2 Symbolic Spatial Relation Mapping
This mapping is informed by prior work on relational scene representation [58, 59], which demonstrates that spatial relations such as above, next to, and on top of align with how people describe and interpret visual layouts. For our system, we adopt the following spatial relation types:
@@ 3.3 Semantic Asset Retrieval with GameTileNet
To align narrative objects with appropriate visual tiles, we adopt a semantic embedding–based retrieval strategy grounded in the GameTileNet dataset [1]. Each tile in the dataset is annotated with structured metadata, including object name, group label, supercategory, and affordance type. These attributes are embedded using the all-MiniLM-L6-v2 Sentence Transformer to construct a searchable index. Narrative objects are encoded using the same model, and tile matches are retrieved via cosine similarity.
@@
GameTileNet classifies tiles into five affordance types adapted from the Video Game Description Language (VGDL) [60]:
@@ 3.6.3 Relation to Hierarchical Narrative Models
Our construction is inspired by prior work on hierarchical symbolic representation for visual narratives [56]. In that model, events are represented at multiple levels, from panel to event to macro-event, and integrated using graph structures that encode semantic, temporal, and multimodal relations.
Although our current system focuses on symbolic graphs from text and image grounding, its scene-level KG resembles the event segment layer in [56]. Both:
@@ 3.7 Visual Rendering and Output
After semantic matching and spatial placement, each tile-based scene is rendered using matched 2D sprite assets. Scenes are organized into multi-layer matrices representing different object types (terrain, environment, interactive objects, items, characters), which are composited in semantic order to preserve depth and spatial logic. Object images are resized (e.g., $1.5\times$), centered within their tiles, and pasted layer by layer to construct the final image. Figure layers are rendered from background to foreground based on their affordance class. We also retain each layer’s numerical matrix for downstream tasks such as symbolic reasoning or gameplay simulation.
The rendering procedure is outlined in Algorithm 3.
@@ 4 Evaluation
We evaluated our narrative-to-scene generation system using 10 LLM-generated stories. Each story is segmented into three key time frames, resulting in a total of 30 scene visualizations and their associated symbolic representations. We assess the quality of the generated outputs from multiple perspectives: tile–object semantic alignment, affordance-layer placement correctness, spatial predicate satisfaction, and qualitative renderings. Optional analysis of temporal consistency and human evaluation is discussed at the end.

### v2

@@ Abstract
Recent advances in large language models (LLMs) enable compelling story generation, but connecting narrative text to playable visual environments remains an open challenge in procedural content generation (PCG). We present a lightweight pipeline that transforms short narrative prompts into a sequence of 2D tile-based game scenes, reflecting the temporal structure of stories. Given an LLM-generated narrative, our system identifies three key time frames, extracts spatial predicates in the form of "Object-Relation-Object" triples, and retrieves visual assets using affordance-aware semantic embeddings from the GameTileNet dataset [chen2025gametilenet]. A layered terrain is generated using Cellular Automata, and objects are placed using spatial rules grounded in the predicate structure. We evaluated our system in ten diverse stories, analyzing tile–object matching, affordance–layer alignment, and spatial constraint satisfaction across frames. This prototype offers a scalable approach to narrative-driven scene generation and lays the foundation for future work on multi-frame continuity, symbolic tracking, and multi-agent coordination in story-centered PCG.
@@ 1 Introduction
Large language models (LLMs) can generate rich, coherent narratives from minimal input, creating new opportunities for procedural content generation (PCG) in games that emphasize narrative and visual storytelling. Yet while narrative generation has advanced rapidly, translating stories into structured, spatially grounded game scenes remains an open challenge. Most PCG systems still produce isolated levels or static layouts without modeling temporal structure and continuity—core aspects of narrative experiences that connect space, time, and character action.
Narratives unfold through sequences of events that shape spatial progression and thematic coherence [mccloud1993understanding, cohn2013visual]. Building on this idea, we present a method for generating sequences of visual game scenes from short narratives. Each story is segmented into three temporal frames (beginning, middle, end) following narrative-structure conventions such as Freytag’s pyramid, capturing key transitions while keeping generation tractable. This segmentation supports partial modeling of story progression and provides a foundation for studying temporal consistency and dynamic storytelling.
Our approach introduces a symbolic-to-visual pipeline that extracts spatial predicates from each frame, maps them to assets in the GameTileNet dataset [chen2025gametilenet], and arranges them on procedurally generated terrains. Object–relation–object triples are derived through LLM prompting and grounded in visual tiles using semantic and affordance-based filtering. Terrain layers are produced with Cellular Automata to ensure connectivity, while objects are placed through rule-based spatial refinement that satisfies adjacency and containment constraints. The resulting layered 2D scenes reflect the structure of the source narrative while maintaining spatial logic relevant to gameplay.
Our contribution lies not in proposing new generative algorithms but in integrating established PCG components, such as Cellular Automata and semantic matching, into a unified narrative-to-scene pipeline emphasizing temporal segmentation and affordance grounding. This coordination supports cross-frame continuity and balances semantic fidelity with affordance validity, a combination rarely explored in prior PCG work. LLM-generated stories serve only as a reproducible testbed for open-vocabulary narrative input; the pipeline itself is agnostic to text source and can process any authored story.
Table 1: Example of generating a tile-based game scene from narrative descriptions. The system extracts spatial relations and semantic objects (left), selects corresponding tile images from the GameTileNet dataset (middle), and renders a visual scene that satisfies both semantic and spatial constraints (right). Top: labeled object layout. Bottom: rendered tile-mapped scene.
@@
We evaluate the system on ten narrative examples, each divided into three frames. Evaluation considers semantic-tile matching, affordance alignment, and spatial-relation satisfaction in the generated scenes. Even with lightweight symbolic rules and open-ended input, the results exhibit coherent spatial organization that reflects narrative intent.
Contributions.
  • A unified pipeline that decomposes short narratives into temporally segmented, layered 2D scenes.
  • A predicate-extraction and semantic-matching process that grounds narrative objects in tile-based assets using affordance-aware filtering.
  • A multi-layer scene synthesis method combining Cellular Automata terrain generation with rule-based spatial placement.
  • An empirical analysis of ten narrative examples demonstrating semantic alignment, affordance coherence, and spatial-relation satisfaction across frames.
The proposed system offers a modular foundation for narrative-to-scene generation, emphasizing temporal segmentation, affordance grounding, and semantic alignment. Although the current prototype does not model agent behavior or dynamic state transitions, it establishes a framework extendable to multi-agent coordination and gameplay-aware narrative PCG. Recent work shows that LLMs can segment narrative events in ways that align with human perception [michelmann2025large], supporting our choice to use structured time frames as the basis for visual scene construction. Table 1 illustrates one such example, where spatial constraints and tile affordances jointly produce semantically coherent game scenes from narrative descriptions.
@@ 2.1 Narrative Theory and Structure in Games
Narrative structure in games has been studied through formalist, experiential, and emergent perspectives. Early work examined the tension between linear storytelling and player agency [aarseth2012narrative, juul2005games]. Later models addressed experiential modes [calleja2009experiential] and story architectures that accommodate agency [domsch2013storyplaying]. Other work connected narrative to spatial layout and symbolic environments [domsch2019space, lindley2005story], while surveys summarize recurring patterns in interactive media [koenitz2024narrative]. These theories provide foundations for incorporating narrative intent into generative pipelines.
@@ 2.2 Narrative-Guided Procedural Content Generation
Procedural content generation (PCG) increasingly leverages narrative structure. Quest and story grammars guide goal-oriented generation [howard2022quests], and language models now produce worlds and events [nasir2024word2world, todd2024gavel]. Emotional arcs and planning constraints improve coherence and engagement [miller2019stories]. PCG methods span grammars, search-based strategies, and PCGML [summerville2018procedural, yannakakis2025procedural], with reinforcement learning extending control to narrative form [togelius2010search]. Layered generation has also appeared in visual storytelling and comics [chen2024collaborative, chen2023customizable], demonstrating ways to link symbolic story input with generative models. Prior text-to-scene and PCG systems generate single scenes or levels [riedl2010narrative, guzdial2018combinatorial, summerville2018procedural] but rarely model temporal continuity or explicit narrative structure. Our work addresses this gap by connecting narrative segmentation with spatial and affordance reasoning, bridging story progression and playable spatial layouts.
@@ 2.3 Tile-Based Generation and Semantic Affordances
Tile-based abstractions remain central to scalable generation. Representative methods include constrained layout via SMT solvers [whitehead2020spatial], evolutionary search [petrovas2022procedural], and grammar-based dungeon layouts [jiang2021synthesizing]. Neural approaches introduce embeddings for generalized level generation [jadhav2021tile] and segmentation for context-aware tilemaps [gabriel2023semantic]. Affordance modeling distinguishes environmental, interactive, and collectible roles [hunicke2004mda], while corpora enable affordance-aware generation [bentley2019videogame]. GameTileNet contributes a dataset of low-resolution tiles annotated with layered affordances [chen2025gametilenet].
@@ 2.4 Semantic Matching and LLM-Guided Grounding
LLMs are increasingly used for scene synthesis and grounding. Prompting and multimodal alignment allow them to act as planners for layouts [volum2022craft, zhou2024scenex]. Studies show that structured input such as scene semantics improves alignment [cao2023comprehensive]. Visual-semantic reasoning frameworks for matching and embedding inform techniques for aligning text predicates with tiles [li2019visual, li2022image]. Cognitive studies suggest LLMs approximate human-like segmentation of narrative events, underscoring their potential for symbolic-neural integration [michelmann2025large].
@@ 2.5 Symbolic Structures in Visual Storytelling
Symbolic and graph-based representations provide structure for storytelling. Knowledge-enhanced generation has been modeled with narrative graphs or multimodal scene graphs [hsu2020knowledge, xu2021imagine, mishra2019storytelling]. Other frameworks explore relational encodings to ensure continuity and causality in stories and games [rivera2024story, akimoto2017computational]. Recent work highlights hierarchical narrative graphs as interpretable intermediaries for multimodal alignment [chen2025hierarchical]. These approaches underscore the importance of symbolic scaffolds for controllable and interpretable storytelling pipelines.
@@ 3 Method: Narrative-to-Scene Generation Pipeline
We propose a structured pipeline that transforms narrative text into visually grounded, tile-based game scenes. As illustrated in Figure 1, the pipeline integrates large language model (LLM)–based story generation with a sequence of symbolic and visual reasoning modules. The process begins with narrative prompting and temporal abstraction, where the story is segmented into key time frames and represented as predicate-style triples. These structured descriptions serve as input for semantic reasoning and symbolic grounding. Although the LLM provides the narrative input, this step serves only to ensure reproducible open-vocabulary test cases; the pipeline itself is model-agnostic and can process any authored story with minimal preprocessing.
Each scene is synthesized through a layered generation process: (1) terrain generation using Cellular Automata ensures navigable base regions; (2) semantic object matching retrieves tile assets aligned with narrative entities based on name, category, and affordance embeddings; and (3) spatial placement with rule-based refinement positions entities according to annotated spatial relations. The final output consists of rendered tile-based scenes and underlying semantic maps, enabling both human-readable visualization and downstream interactive applications.
Figure 1: Pipeline for generating a tile-based game scene from an LLM-generated narrative. Modules include narrative parsing, semantic matching, procedural terrain generation, and layered scene rendering.
@@ 3.1 Narrative Prompting and Frame Extraction
Inspired by Word2World [nasir2024word2world], we use a large language model to generate a short adventure story and extract three key narrative moments. Each moment is structured as symbolic predicates describing relationships between story elements, represented as [Object] [Relation] [Object] triples. These structured predicates form the basis for downstream spatial reasoning and visual scene generation.
The process uses two prompts:
  • Prompt_1: “Generate a short adventure story (about 100 words).”
  • Prompt_2: “Extract three key time frames and describe each with [Object] [Relation] [Object] triplets.”
@@
Table 2 shows the extracted time frames, the corresponding predicate triples, and the tile-based scenes rendered by our pipeline.
@@
Hollow oak contains ancient map Elara stands near hollow oak Sunlight filters through forest canopy Elara walks along rocky path Wild creatures hide behind dense bushes Treacherous paths lead to Crystal Cavern Crystal Cavern entrance glows with shimmering light Guardian dragon sits atop crystal throne Elara stands before Guardian dragon
@@ 3.2 Symbolic Spatial Relation Mapping
This mapping is informed by prior work on relational scene representation [krishna2017visual, johnson2017clevr], which demonstrates that spatial relations such as above, next to, and on top of align with how people describe and interpret visual layouts. For our system, we adopt the following spatial relation types:
@@ 3.3 Semantic Asset Retrieval with GameTileNet
To align narrative objects with appropriate visual tiles, we adopt a semantic embedding–based retrieval strategy grounded in the GameTileNet dataset [chen2025gametilenet]. Each tile in the dataset is annotated with structured metadata, including object name, group label, supercategory, and affordance type. These attributes are embedded using the all-MiniLM-L6-v2 Sentence Transformer to construct a searchable index. Narrative objects are encoded using the same model, and tile matches are retrieved via cosine similarity.
@@
GameTileNet classifies tiles into five affordance types adapted from the Video Game Description Language (VGDL) [schaul2013video]:
@@ 3.6.3 Relation to Hierarchical Narrative Models
Our construction is inspired by prior work on hierarchical symbolic representation for visual narratives [chen2025hierarchical]. In that model, events are represented at multiple levels, from panel to event to macro-event, and integrated using graph structures that encode semantic, temporal, and multimodal relations.
Although our current system focuses on symbolic graphs from text and image grounding, its scene-level KG resembles the event segment layer in [chen2025hierarchical]. Both:
@@ 3.7 Visual Rendering and Output
After semantic matching and spatial placement, each tile-based scene is rendered using matched 2D sprite assets. Scenes are organized into multi-layer matrices representing different object types (terrain, environment, interactive objects, items, characters), which are composited in semantic order to preserve depth and spatial logic. Object images are resized (e.g., $1.5\times$), centered within their tiles, and pasted layer by layer to construct the final image. Figure layers are rendered from background to foreground based on their affordance class. We also retain each layer’s numerical matrix for downstream tasks such as symbolic reasoning or gameplay simulation. The rendering procedure is outlined in Algorithm 3.
@@ 4 Evaluation
We evaluated our narrative-to-scene generation system using 10 LLM-generated stories. Each story is segmented into three key time frames, resulting in a total of 30 scene visualizations and their associated symbolic representations. We assess the quality of the generated outputs from multiple perspectives: tile–object semantic alignment, affordance-layer placement correctness, spatial predicate satisfaction, and qualitative renderings.